# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 3.4.2. Smoothing of $\mathcal{X}_{0}$
Recall that the affinization map $f_{0}:\mathcal{X}_{0}\rightarrow\overline{\mathcal{X}}_{0}$ is a projective surjective birational morphism. Since $\overline{\mathcal{X}}_{0}$ is affine, then, by [51, Proposition III.8.5] (a degenerate case of the Leray spectral sequence), we have
$H^{2}(\mathcal{X}_{0},T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}})=H^{0}(\overline{\mathcal{X}}_{0},R^{2}(f_{0})_{\star}\,T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}})\,.$
Fibers of $f_{0}$ over points in $\overline{\mathcal{X}}_{0}$ distinct from the central point $x_{0}$ are of dimension at most one, and so, by [51, Corollary III.11.2], the sheaf $R^{2}(f_{0})_{\star}\,T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}$ is supported on the central point $x_{0}$. Hence, it is enough to show that its stalk at $x_{0}$ is 00:
$\left(R^{2}(f_{0})_{\star}\,T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}\right)_{x_{0}}=0\,.$ (3.8)
As the sheaf $R^{2}(f_{0})_{\star}\,T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}$ is supported $x_{0}$, the formal completion of the stalk at $x_{0}$ is equal to the stalk at $x_{0}$ itself, and so, by the theorem on formal functions [51, Theorem 11.1], we have
[…]
where for every $k\in\mathbb{Z}_{\geq 0}$, $\mathcal{X}_{0}^{(k)}$ is the projective subscheme of $\mathcal{X}_{0}$ obtained by taking the scheme-theoretic preimage by $f_{0}$ of the $k$-th infinitesimal neighborhood of $x_{0}$ in $\overline{\mathcal{X}}_{0}$. In particular, for every $k\in\mathbb{Z}_{\geq 0}$, the reduced underlying scheme of $\mathcal{X}_{0}^{(k)}$ is the fiber $\mathcal{X}_{0}^{(0)}=f_{0}^{-1}(x_{0})$. As all proper 2-dimensional irreducible of components of $\mathcal{X}_{0}$ are contained in $\mathcal{X}_{0}^{(0)}$, it follows that, for every $k\in\mathbb{Z}_{\geq 0}$, the schemes $\mathcal{X}_{0}^{(k)}$ and $\mathcal{X}_{0}^{(0)}$ only differ along the finitely many curves where $\mathcal{X}_{0}^{(0)}$ intersect the non-proper irreducible components of $\mathcal{X}_{0}$. In particular, the kernel and cokernel of the restriction map
$T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{X}_{0}^{(k)}}\longrightarrow T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{X}_{0}^{(0)}}$
are supported in dimension $1$, so have vanishing $H^{2}$, and so the corresponding long exact sequences in cohomology imply that, for every $k\in\mathbb{Z}_{\geq 0}$, we have
$H^{2}(\mathcal{X}_{0}^{(k)},T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{X}_{0}^{(k)}})=H^{2}(\mathcal{X}_{0}^{(0)},T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{X}_{0}^{(0)}})\,.$
Thus, in order to prove the vanishing in Equation (3.8), and so Lemma 3.25, it is enough to show that
$H^{2}(\mathcal{X}_{0}^{(0)},T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{X}_{0}^{(0)}})=0\,.$ (3.9)
The irreducible components of $\mathcal{X}_{0}^{(0)}$ are the proper surfaces $X^{v}_{\mathscr{P}}$, indexed by the interior vertices $v\in\mathscr{P}^{[0]}$, and the proper curves $X^{e}_{\mathscr{P}}$ indexed by the set $\mathscr{P}^{[1]}_{\mathcal{E}}$ of interior edges $e\in\mathscr{P}^{[1]}$ connecting pairs of boundary vertices. We denote by $\mathcal{X}^{(0)}_{0}=\mathcal{Z}\cup\mathcal{E}$ the corresponding decomposition, where
$\mathcal{Z}:=\bigcup_{\begin{subarray}{c}v\in\mathscr{P}^{[0]}\\ v\notin\partial P\end{subarray}}X^{v}\,\,\,\,\,\,\text{and}\,\,\,\,\,\,\mathcal{E}:=\bigcup_{\begin{subarray}{c}v\in\mathscr{P}^{[1]}_{\mathcal{E}}\end{subarray}}X^{e}\,.$
Note that the intersection $\mathcal{Z}\cap\mathcal{E}$ is a finite union of triple points of $\mathcal{X}_{0}$. Tensoring the short exact sequence
$0\rightarrow\mathcal{O}_{\mathcal{X}^{(0)}_{0}}\rightarrow\mathcal{O}_{\mathcal{Z}}\oplus\mathcal{O}_{\mathcal{E}}\rightarrow\mathcal{O}_{\mathcal{Z}\cap\mathcal{E}}\rightarrow 0$
with the locally free log tangent bundle $T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}$ produces a short exact sequence
@@
$0\rightarrow T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{X}^{(0)}_{0}}\rightarrow T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{Z}}\oplus T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{E}}\rightarrow T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{Z}\cap\mathcal{E}}\rightarrow 0\,.$
Using the vanishing of the $H^{1}$ and $H^{2}$ groups implied by $\dim\mathcal{E}=1$ and $\dim\mathcal{Z}\cap\mathcal{E}=0$ in the corresponding long exact sequence in cohomology, we obtain
$H^{2}(\mathcal{X}_{0}^{(0)},T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{X}_{0}^{(0)}})=H^{2}(\mathcal{Z},T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{Z}})\,.$ (3.10)
By Lemma 3.8, the dualizing sheaf $\omega_{\mathscr{X}_{0}}$ is trivial, and so, by adjunction, the dualizing sheaf $\omega_{\mathcal{Z}}$ of $\mathcal{Z}$ is given by the line bundle $\omega_{\mathcal{Z}}=\mathcal{O}_{\mathcal{Z}}(-\partial\mathcal{Z})$. Hence, by Serre duality on $\mathcal{Z}$, and using that $T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{Z}}$ is locally free on $\mathcal{Z}$, we obtain
$H^{2}(\mathcal{Z},T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{Z}})=H^{0}(\mathcal{Z},(\Omega_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{Z}})(-\partial\mathcal{Z}))^{\vee}\,.$ (3.11)
By Lemma 3.24, we have $H^{0}(\mathcal{Z},(\Omega_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\,\mathbb{C}^{\dagger}}|_{\mathcal{Z}})(-\partial\mathcal{Z}))=0$, and so we deduce using Equations (3.10)-(3.11) that the vanishing of Equation (3.9) holds, and this concludes the proof of Lemma 3.25. ∎
Finally, we are ready to state the existence of a smoothing of $\mathcal{X}_{0}$.
@@
    The total space $\mathcal{X}$ is smooth.
@@
    The central fiber $\mathcal{X}_{0}$ is a normal crossing divisor in $\mathcal{X}_{\mathscr{P}}$.
@@
    ${\pi_{0}}^{\dagger}\colon{\mathcal{X}_{0}}^{\dagger}=(\mathcal{X}_{0},\mathcal{M}_{\mathcal{X}_{0}})\longrightarrow\mathrm{Spec}\,\mathbb{C}^{\dagger}$
@@
Since the obstruction group $H^{2}(\mathcal{X}_{0},T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\mathbb{C}^{\dagger}})$ vanishes by Lemma 3.25, the result follows by log smooth deformation theory as in [62, Corollary 2.4]. ∎
@@ 3.5. Open Kulikov degenerations
We define below open Kulikov degenerations as “semistable smoothings” of a generic open Kulikov surfaces.
@@
An open Kulikov degeneration is a proper morphism $\pi:\mathcal{X}\rightarrow\Delta:=\mathrm{Spf}\,\mathbb{C}[\![t]\!]$ of formal schemes satisfying the following conditions.
@@
    The central fiber $\mathcal{X}_{0}$ is a generic open Kulikov surface – as in Definition 3.19.
@@
By definition, the central fiber of an open Kulikov degeneration is a $d$-semistable generic open Kulikov surface. Conversely, by Theorem 3.26, every $d$-semistable generic open Kulikov surface is the central fiber of an open Kulikov degeneration.
Recall from §3.2 that every open Kulikov surface $\mathcal{X}_{0}$ has an affinization $\overline{\mathcal{X}}_{0}$. The following result describes the affinization of open Kulikov degenerations.
@@
Let $\pi\colon\mathcal{X}\to\Delta$ be a semistable smoothing of a generic open Kulikov surface $\mathcal{X}_{0}$. Then, the affinization $\overline{\mathcal{X}}\coloneqq\mathrm{Spec}\,H^{0}(\mathcal{X},\mathcal{O}_{\mathcal{X}})$ is a smoothing of $\overline{\mathcal{X}}_{0}\coloneqq\mathrm{Spec}\,H^{0}(\overline{\mathcal{X}}_{0},\mathcal{O}_{\overline{\mathcal{X}}_{0}})$, and has Gorenstein canonical singularities. Moreover, $\mathcal{X}$ has trivial canonical line bundle, and the affinization morphism $f\colon\mathcal{X}\to\overline{\mathcal{X}}$ is a crepant resolution of singularities.
@@
By Lemma 3.10, we have $H^{1}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, and so it follows from [96, Theorem 1.4] (see also [26, Theorem 3.1]) that $f$ is a flat deformation of $f_{0}$. To prove that $\overline{\mathcal{X}}$ has Gorenstein canonical singularities, first note that $\mathcal{X}$ has trivial canonical line bundle – this can be shown analogously as in [33, Theorem 5.10, 3)], where a similar result is proved in the context of degenerations of K3 surfaces. Hence, since $\mathcal{X}$ is a smooth threefold with trivial canonical class, by [60, Corollary 1.5] it follows that $\overline{\mathcal{X}}$ has Gorenstein canonical singularities. Note that [60, Corollary 1.5] is stated in the context of projective morphism between normal algebraic varieties. In this context, the result follows from the base point free theorem [60, Theorem 1.3]. The same argument shows that the result for projective morphisms of formal schemes follows from the corresponding version of the base point free theorem in [75, §11].
It remains to show that $\overline{\mathcal{X}}\rightarrow\Delta$ is a smoothing of $\mathcal{X}_{0}$. It is enough to show that the rigid-analytic generic fiber $\overline{\mathcal{X}}_{\eta}$ of $\overline{\mathcal{X}}$ is smooth. Denote by $\mathcal{X}_{\eta}$ the rigid-analytic generic fiber of $\mathcal{X}$, which is smooth since $\mathcal{X}\rightarrow\Delta$ is smoothing. Standard results on birational geometry of surfaces generalize to rigid-analytic surfaces by [95, §5]. Hence, it follows from the triviality of the canonical class of $\mathcal{X}_{\eta}$ has trivial that the birational map $\mathcal{X}_{\eta}\rightarrow\overline{\mathcal{X}}_{\eta}$ can only contract $(-2)$-curves. Hence, it is enough to show that $\mathcal{X}_{\eta}$ does not contain any $(-2)$-curve. If $C_{\eta}$ were a $(-2)$-curve in $\mathcal{X}_{\eta}$, then up to replacing $\mathcal{X}$ by a log resolution $\widetilde{\mathcal{X}}$ of a base change, $C_{\eta}$ would admit a model with central fiber a curve $C_{0}$ in $\widetilde{\mathcal{X}}_{0}$ containing no 0-dimensional stratum of $\widetilde{\mathcal{X}}_{0}$. Such a curve $C_{0}$ would necessarily be either an internal $(-2)$-curve or a degenerate $(-2)$-curve as in Definition 3.17, in contradiction with the assumption that the open Kulikov surface $\mathcal{X}_{0}$ is generic as in Definition 3.19. ∎
Remark 3.30.
Since $\overline{\mathcal{X}}$ is an affine formal smoothing, it follows from Elkik’s theorem [29, Theorem 7, p582] – see also [92, Proposition 3.3.1], that $\overline{\mathcal{X}}$ is algebraizable over $\mathrm{Spec}\,\mathbb{C}[\![t]\!]$. Moreover, $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is a formal blow-up, and so is algebraizable in the category of algebraic spaces over $\mathrm{Spec}\,\mathbb{C}[\![t]\!]$ by [19, Theorem 3.2]. Since $H^{2}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, and $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is projective, it follows that $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is in fact algebraizable in the category of schemes over $\mathrm{Spec}\,\mathbb{C}[\![t]\!]$.
@@
  • (iii)
    The affinization morphism $f:\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is projective.
@@
The only non-immediate implication is i) implies ii). By Lemma 3.10, we have $H^{2}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, so there are no obstructions to formal deformations over $\Delta$ of an ample line bundle on $\mathcal{X}_{0}$, and so the result follows. ∎

### v3

@@ 3.3.2. Smoothing of $\mathcal{X}_{0}$
By Lemma 3.8, the dualizing sheaf $\omega_{\mathcal{X}_{0}}$ of $\mathcal{X}_{0}$ is trivial, and so, by Serre duality, $H^{2}(\mathcal{X}_{0},T_{\mathcal{X}_{0}^{\dagger}/\mathrm{pt}_{\mathbb{N}}})$ is dual to $H^{0}_{c}(\mathcal{X}_{0},\Omega_{\mathcal{X}_{0}^{\dagger}/\mathrm{pt}_{\mathbb{N}}})$, where $H^{i}_{c}$ denotes compactly supported coherent cohomology groups – see [55] or [93, Tag 0G59], and $\Omega_{\mathcal{X}_{0}^{\dagger}/\mathrm{pt}_{\mathbb{N}}}$ is the log cotangent bundle of $\mathcal{X}_{0}^{\dagger}$ over the standard log point $\mathrm{pt}_{\mathbb{N}}$. Hence, to prove Lemma 3.18, it suffices to show that $H^{0}_{c}(\mathcal{X}_{0},\Omega_{\mathcal{X}_{0}^{\dagger}/\mathrm{pt}_{\mathbb{N}}})=0$.
To do this, we adapt to our situation the argument given in the proof of [35, Lemma 5.9] (where the log cotangent bundle is denoted by $\Lambda^{1}_{X}$) in the context of degenerations of K3 surfaces – see also [88, Theorem 5.2] for a similar result in the context of degenerations of Enriques surfaces. We denote by $(P,\mathscr{P})$ the dual intersection complex of $\mathcal{X}_{0}$, with vertices $v\in\mathscr{P}^{[0]}$ (resp. edges $e\in\mathscr{P}^{[1]}$ and triangular faces $f\in\mathscr{P}^{[2]}$) in one-to-one correspondence with the irreducible components $X^{v}$ (resp. double curves $X^{e}$ and triple points $X^{f}$) of $\mathcal{X}_{0}$. By [68, Proof of Theorem 4.2] (see also [35, Corollary 3.6] or [33, Proof of Lemma 2.21] where Equation (3.6) below is taken as the starting point of the discussion), it follows from the definition of the log cotangent bundle that we have a short exact sequence
$0\rightarrow\Omega_{\mathcal{X}_{0}}/\tau\rightarrow\Omega_{\mathcal{X}_{0}^{\dagger}/\mathrm{pt}_{\mathbb{N}}}\rightarrow F\rightarrow 0\,$, (3.4)
where $\Omega_{\mathcal{X}_{0}}$ is the usual cotangent sheaf of $\mathcal{X}_{0}$, $\tau$ is the torsion subsheaf of $\Omega_{\mathcal{X}_{0}}$, and the cokernel $F$ sits via the Poincaré residue map into the short exact sequence
$0\rightarrow\mathcal{O}_{\mathcal{X}_{0}}\rightarrow\bigoplus_{v\in\mathscr{P}^{[0]}}\mathcal{O}_{X^{v}}\rightarrow F\rightarrow 0\,.$
The long exact sequence in cohomology with compact support associated to the short exact sequence (3.4) is of the form
@@
$0\rightarrow H^{0}_{c}(\mathcal{X}_{0},\Omega_{\mathcal{X}_{0}}/\tau)\rightarrow H^{0}_{c}(\mathcal{X}_{0},\Omega_{\mathcal{X}_{0}^{\dagger}/\mathrm{pt}_{\mathbb{N}}})\rightarrow H^{0}_{c}(\mathcal{X}_{0},F)\xrightarrow{\epsilon}H^{1}_{c}(\mathcal{X}_{0},\Omega_{\mathcal{X}_{0}}/\tau)\,.$ (3.5)
On the other hand, $\tau$ is exactly the kernel of the natural restriction map $\Omega_{\mathcal{X}_{0}}\rightarrow\oplus_{v\in\mathscr{P}^{[1]}}\Omega_{X^{v}}$ (see [35, p77]), so $\Omega_{\mathcal{X}_{0}}/\tau$ is a subsheaf of $\oplus_{v\in\mathscr{P}^{[1]}}\Omega_{X^{v}}$. Every proper irreducible component $X^{v}$ of $\mathcal{X}_{0}$ is a proper log Calabi–Yau surface, satisfying $H^{0}(X^{v},\Omega_{X^{v}})\subset H^{1}(X^{v},\mathbb{C})$ by the Hodge decomposition, and so $H^{0}(X^{v},\Omega_{X^{v}})=0$ since $H^{1}(X^{v},\mathbb{C})=0$. Thus, we obtain $H^{0}_{c}(\mathcal{X}_{0},\Omega_{\mathcal{X}_{0}}/\tau)=0$, and so it remains to prove that the map $\epsilon$ in (3.5) is injective.
It follows from the exact sequence
$0\rightarrow\mathcal{O}_{\mathcal{X}_{0}}\rightarrow\bigoplus_{v\in\mathscr{P}^{[0]}}\mathcal{O}_{X^{v}}\rightarrow\bigoplus_{e\in\mathscr{P}^{[1]}}\mathcal{O}_{X^{e}}\rightarrow\bigoplus_{f\in\mathscr{P}^{[2]}}\mathcal{O}_{X^{f}}\rightarrow 0\,$,
that we have
$F=\mathrm{Ker}\left(\bigoplus_{e\in\mathscr{P}^{[1]}}\mathcal{O}_{X^{e}}\rightarrow\bigoplus_{f\in\mathscr{P}^{[2]}}\mathcal{O}_{X^{f}}\right)\,$, (3.6)
and so
$H^{0}_{c}(\mathcal{X}_{0},F)=\mathrm{Ker}\left(\bigoplus_{e\in\mathscr{P}^{[1]}}H^{0}_{c}(X^{e},\mathcal{O}_{X^{e}})\rightarrow\bigoplus_{f\in\mathscr{P}^{[2]}}H^{0}_{c}(X^{f},\mathcal{O}_{X^{f}})\right)\,.$
Since $H^{0}_{c}(X^{e},\mathcal{O}_{X^{e}})=0$ if $X^{e}$ is non-compact, that is, $e\subset\partial P$, and $H^{0}_{c}(X^{e},\mathcal{O}_{X^{e}})=\mathbb{C}$ otherwise, $H^{0}_{c}(\mathcal{X}_{0},F)$ consists of tuples $(a_{e})_{e\in\mathscr{P}^{[1]}}\in\mathbb{C}^{\mathscr{P}^{[1]}}$, such that $a_{e}=0$ if $e\subset\partial P$, and such that $a_{e_{1}}+a_{e_{2}}+a_{e_{3}}=0$ if $e_{1}$, $e_{2}$, $e_{3}$ are three edges bounding a triangle $f\in\mathscr{P}^{[2]}$.
On the other hand, for every $e\in\mathscr{P}^{[1]}$ such that the double curve $X^{e}$ is proper, $X^{e}$ is a rational curve, and so $H^{0}(X^{e},\Omega_{X^{e}})=0$. Hence, the short exact sequence
$0\rightarrow\Omega_{\mathcal{X}_{0}}/\tau\rightarrow\bigoplus_{v\in\mathscr{P}^{[0]}}\Omega_{X^{v}}\rightarrow\bigoplus_{e\in\mathscr{P}^{[1]}}\Omega_{X^{e}}\rightarrow 0$
implies that
$H^{1}_{c}(\mathcal{X}_{0},\Omega_{\mathcal{X}_{0}}/\tau)=\mathrm{Ker}\left(\bigoplus_{v\in\mathscr{P}^{[0]}}H^{1}_{c}(X^{v},\Omega_{X^{v}})\rightarrow\bigoplus_{e\in\mathscr{P}^{[1]}}H^{1}_{c}(X^{e},\Omega_{X^{e}})\right)\,.$
As in the proof of [35, Lemma 5.9], the map $\epsilon$ in (3.5) is the restriction of the map
$\delta:\bigoplus_{e\in\mathscr{P}^{[1]}}H^{0}_{c}(X^{e},\mathcal{O}_{X^{e}})$ $\longrightarrow\bigoplus_{v\in\mathscr{P}^{[0]}}H^{1}_{c}(X^{v},\Omega_{X^{v}})$ (3.7)
$(a_{e})_{e\in\mathscr{P}^{[1]}}$ $\longrightarrow\left(\sum_{e\in\mathscr{P}^{[1]}}a_{e}[X^{e}]_{v}\right)_{v\in\mathscr{P}^{[0]}}\,$,
where $X^{e}]_{v}$ is the class of the compact curve $X^{e}$ in the surface $X^{v}$ if $X^{e}\subset X^{v}$, that is, if $e$ is an edge incident to the vertex $v$, and $X^{e}]_{v}=0$ else. We can now prove the injectivity of $\epsilon$. Let $(a_{e})_{e\in\mathscr{P}^{[1]}}\in H^{0}_{c}(\mathcal{X}_{0},F)$ such that $\delta((a_{e})_{e})=0$. Since the map $\epsilon$ is the restriction of the map $\delta$ given by (3.7), we have $\sum_{e\in\mathscr{P}^{[1]}}a_{e}[X^{e}]_{v}=0$ in $H^{2}_{c}(X^{v},\mathbb{C})$ for every $v\in\mathscr{P}^{[0]}$. For every vertex $v\in\mathscr{P}^{[0]}$, define the distance to the boundary $d(v)$ as $d(v)=0$ if $v\in\partial P$, and else as the smallest integer such that there exists a sequence $(e_{1},\dots,e_{d(v)})$ of edges $e_{i}\in\mathscr{P}^{[1]}$ such that $e_{1}$ is adjacent to $v$, $e_{i+1}$ is adjacent to $e_{i}$, and $e_{d(v)}$ is adjacent to a vertex in $\partial P$. We will prove that $a_{e}=0$ for all $e\in\mathscr{P}^{[1]}$, by showing by induction on $d(v)\geq 0$ that, for every vertex $v\in\mathscr{P}^{[0]}$, we have $a_{e}=0$ for all edges $e\in\mathscr{P}^{[1]}$ adjacent to $v$.
For the initialization of the induction, note that if $v\in\mathscr{P}^{[0]}$ is such that $d(v)=0$, then $v\in\partial P$. Denote by $e_{0},e_{1},\cdots,e_{n}$ the edges adjacent to $v$, cyclically ordered so that $e_{0},e_{n}\subset\partial P$. Then, we have $a_{e_{0}}=a_{e_{n}}=0$, and so the linear relation $\sum_{i=1}^{n-1}a_{e_{i}}[X^{e_{i}}]_{v}=0$ in $H^{2}_{c}(X^{v},\mathbb{C})$. By Lemma 3.17, the classes $X^{e_{i}}]_{v}$ for $1\leq i\leq n-1$ are linearly independent in $H^{2}_{c}(X^{v},\mathbb{C})$, and so we conclude that $a_{e_{i}}=0$ for all $1\leq i\leq n-1$.
For the induction step, let $d\in\mathbb{Z}_{\geq 1}$, and assume that the result holds for all vertices $v^{\prime}$ with $d(v^{\prime})<d$. Let $v\in\mathscr{P}^{[0]}$ such that $d(v)=d$. Then, there exists an edge $e_{1}$ adjacent to $v$ which is also adjacent to a vertex $v^{\prime}\in\mathscr{P}^{[0]}$ with $d(v^{\prime})<d$. Let $f\in\mathscr{P}^{[2]}$ be a triangle with side $e_{1}$, and denote by $e_{2}$ the other side adjacent to $v$, and by $e_{3}$ the other side adjacent to $v^{\prime}$. By the induction hypothesis, we have $a_{e_{1}}=a_{e_{3}}=0$, and so it follows from the condition $a_{e_{1}}+a_{e_{2}}+a_{e_{3}}=0$ that we also have $a_{e_{2}}=0$. Therefore, the linear relation $\sum_{e\in\mathscr{P}^{[1]}}a_{e}[X^{e}]_{v}=0$ only involves boundary divisor of $X^{v}$ distinct from the consecutive components $X^{e_{1}}$ and $X^{e_{2}}$. Hence, by Lemma 3.16, such boundary divisors are linearly independent in $H^{2}(X^{v},\mathbb{C})$, and so $a_{e}=0$ for all edges $e$ adjacent to $v$. This concludes the proof by induction that the map $\epsilon$ in (3.5) is injective, and so the proof of Lemma 3.18.
Finally, we are ready to state the existence of a formal smoothing of $\mathcal{X}_{0}$. Recall that a formal smoothing of $\mathcal{X}_{0}$ is the data of a formal scheme $\mathcal{X}$ and of a rig-smooth morphism of formal schemes $\pi\colon\mathcal{X}\to\Delta=\mathrm{Spf}\,\mathbb{C}\llbracket t\rrbracket$ (i.e. a morphism with smooth rigid-analytic generic fiber [93, Tag 0GCN]), such that $\mathcal{X}_{0}=\pi^{-1}(0)$.
@@
    The total space $\mathcal{X}$ is regular.
@@
    ${\pi_{0}}^{\dagger}\colon{\mathcal{X}_{0}}^{\dagger}=(\mathcal{X}_{0},\mathcal{M}_{\mathcal{X}_{0}})\longrightarrow\mathrm{pt}_{\mathbb{N}}$
@@
Since the obstruction group $H^{2}(\mathcal{X}_{0},T_{\mathcal{X}_{0}^{\dagger}/\mathrm{Spec}\mathbb{C}^{\dagger}})$ vanishes by Lemma 3.18, the result follows by log smooth deformation theory as in [68, Corollary 2.4]. ∎
@@ 3.4. Open Kulikov degenerations
We define below open Kulikov degenerations as “semistable smoothings” of open Kulikov surfaces.
@@
An open Kulikov degeneration is a formal smoothing $\pi:\mathcal{X}\rightarrow\Delta:=\mathrm{Spf}\,\mathbb{C}[\![t]\!]$ satisfying the following conditions.
@@
    The central fiber $\mathcal{X}_{0}$ is an open Kulikov surface.
@@
By definition, the central fiber of an open Kulikov degeneration is a $d$-semistable open Kulikov surface. Conversely, by Theorem 3.19, every $d$-semistable open Kulikov surface is the central fiber of an open Kulikov degeneration. Recall from §3.2 that every open Kulikov surface $\mathcal{X}_{0}$ has an affinization $\mathcal{X}^{\mathrm{can}}_{0}$. The following result describes the affinization of open Kulikov degenerations.
@@
Let $\pi\colon\mathcal{X}\to\Delta$ be a semistable smoothing of an open Kulikov surface $\mathcal{X}_{0}$. Then, the following holds:
  • i)
    The affinization $\mathcal{X}^{\mathrm{can}}\coloneqq\mathrm{Spec}\,H^{0}(\mathcal{X},\mathcal{O}_{\mathcal{X}})$ is a formal deformation of $\mathcal{X}^{\mathrm{can}}_{0}$, and has Gorenstein canonical singularities.
  • ii)
    The rigid-analytic generic fiber $\mathcal{X}^{\mathrm{can}}_{\eta}$ of $\mathcal{X}^{\mathrm{can}}$ has at worst Du Val (ADE) singularities.
  • iii)
    $\mathcal{X}$ has trivial canonical line bundle, and the affinization morphism $f\colon\mathcal{X}\to\mathcal{X}^{\mathrm{can}}$ is a crepant resolution of singularities.
@@
We first prove Theorem 3.22 i) and iii). By Lemma 3.10, we have $H^{1}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, and so it follows from [100, Theorem 1.4] (see also [27, Theorem 3.1]) that $f$ is a flat deformation of $f_{0}$. To prove that $\mathcal{X}^{\mathrm{can}}$ has Gorenstein canonical singularities, first note that $\mathcal{X}$ has trivial canonical line bundle – this can be shown analogously as in [35, Theorem 5.10, 3)], where a similar result is proved in the context of degenerations of K3 surfaces. Hence, since $\mathcal{X}$ is a smooth threefold with trivial canonical class, it follows by [66, Corollary 1.5] that $\mathcal{X}^{\mathrm{can}}$ has Gorenstein canonical singularities. Note that [66, Corollary 1.5] is stated in the context of projective morphism between normal algebraic varieties. In this context, the result follows from the base point free theorem [66, Theorem 1.3]. The same argument shows that the result for projective morphisms of formal schemes follows from the corresponding version of the base point free theorem in [80, §11].
To prove Theorem 3.22 ii), denote by $\mathcal{X}_{\eta}$ the rigid-analytic generic fiber of $\mathcal{X}$, which is smooth since $\mathcal{X}\rightarrow\Delta$ is smoothing. Standard results on birational geometry of surfaces generalize to rigid-analytic surfaces by [99, §5]. Hence, it follows from the triviality of the canonical class of $\mathcal{X}_{\eta}$ that the birational map $\mathcal{X}_{\eta}\rightarrow\mathcal{X}^{\mathrm{can}}_{\eta}$ can only contract (2)(-2)-curves, and so that $\mathcal{X}^{\mathrm{can}}_{\eta}$ has at worst Du Val (ADE) singularities. ∎
The following result describes the relative Picard group of an open Kulikov degeneration.
Lemma 3.23.
Let $\pi\colon\mathcal{X}\to\Delta$ be a semistable smoothing of an open Kulikov surface $\mathcal{X}_{0}$. Then, the restriction of line bundles to the central fiber induces an isomorphism
$\mathrm{Pic}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\simeq\mathrm{Pic}(\mathcal{X}_{0})\,.$
Proof.
By Lemma 3.10, we have $H^{1}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=H^{2}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, and so, by formal deformation theory of line bundles, we have $\mathrm{Pic}(\mathcal{X})\simeq\mathrm{Pic}(\mathcal{X}_{0})$. On the other hand, it follows from the explicit description of $\mathcal{X}_{0}^{\mathrm{can}}$ given by Lemma 3.14 that $\mathrm{Pic}(\mathcal{X}_{0}^{\mathrm{can}})=0$, $H^{1}(\mathcal{X}_{0}^{\mathrm{can}},\mathcal{O}_{\mathcal{X}_{0}^{\mathrm{can}}})=H^{2}(\mathcal{X}_{0}^{\mathrm{can}},\mathcal{O}_{\mathcal{X}_{0}^{\mathrm{can}}})=0$, and so that $\mathrm{Pic}(\mathcal{X}_{0}^{\mathrm{can}})=0$. ∎
@@
  • iii)
    The affinization morphism $f:\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ is projective.
@@
The only non-immediate implication is i) implies ii). By Lemma 3.23, so there are no obstructions to formal deformations over $\Delta$ of an ample line bundle on $\mathcal{X}_{0}$, and so the result follows. ∎